OpenStudy (anonymous):

Donna deposits 8000 into an account that pays simple interest at a rate of 6% per year. How much interest will she be paid in the first 6 years?

OpenStudy (anonymous):

is the answer 2880?

Parth (parthkohli):

Interest = P * R * T

OpenStudy (anonymous):

jiberish

OpenStudy (anonymous):

whats p r t?

Parth (parthkohli):

\( \color{Black}{\Rightarrow Interest = 8000 \cdot 0.06 \cdot 6}\)

Parth (parthkohli):

P is Principal R is rate T is time